### v1

@@ 4.1.3.2 Infer async transformation
3
4def fetch_data(year):
@@
3
4def test_fetch_data():
@@
$\Downarrow$ Infer async transformation $\Downarrow$
@@
3
4async def fetch_data(year):
@@
Identifying asynced functions For a given file that was migrated in the LLM migration round (Section 4.1.2), PyMigTool builds a abstract syntax tree (AST) of the code before migration and after migration. PyMigTool then finds the asynchronous function definitions in the code after migration. For each such function, PyMigTool finds functions with the same qualified name in the AST of the code before migration. If such a function exists in the AST before migration, but was not asynchronous, we consider the function as asynced. For example, in Figure 8, the function fetch_data is asynced, because it was synchronous before migration, but is asynchronous after migration. If a corresponding function is not found at all, it means that the function is newly added due to the migration, and we do not consider it as asynced, because it is introduced by the LLM and does not have callers in other parts of the project.
@@
Identifying to-async functions and to-await calls If a function a calls a function b, and b calls a function c, then a is said to transitively call b and c. For a given asynced function, PyMigTool finds all its transitive callers using the call graph, and marks them as to-async. For example, in Figure 8, the function fetch_data is asynced, and it is called by the function test_fetch_data, therefore, PyMigTool marks test_fetch_data as to-async. It then again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await. For example, in Figure 8, the call to fetch_data in the function test_fetch_data is marked as to-await, because it is a call to an asynced function. The to-async functions need the async keyword to be added to their signature, and the to-await calls need the await keyword to be added to them.
Applying the changes PyMigTool uses LibCST liblibcst to parse the code, search code elements to be modified, and applies the following changes:
  1. 1.
    If a function definition is marked to-async, and it is not already asynchronous, then PyMigTool adds the async keyword to the function definition, for example the function definition test_fetch_data in Figure 8.
  2. 2.
    If a call is marked to-await, and it is already not awaited, PyMigTool adds the await keyword to the call, for example the call to fetch_data in test_fetch_data in Figure 8.
  3. 3.
    If a test function is asynced or is marked to-async, PyMigTool decorates the test function with pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests. For example, in Figure 8, the function test_fetch_data is decorated with pytest.mark.asyncio.
After finishing the async transfomations, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportasync.
4.2 Evaluating PyMigTool
We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects. By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high test coverage, so that the evaluation is reliable. We run each migration just one time, unlike in Section 3, because we found in that section that the LLM does not produce highly divergent migration results when run multiple times on the code. The setup we use is same as the setup in LABEL:rq:unseen (LABEL:sec:rq-llm-unseen), where we evaluate the LLMs on a set of migrations that were not used in training the LLMs.
We aim to answer the following research questions:
  1. RQ-B1
@@
    How many migrations can PyMigTool perform correctly? We measure how well PyMigTool performs on the simulated migrations without any manual intervention.
  2. RQ-B2
@@
    How much manual effort is required to complete the incomplete migrations? While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention.
4.2.1 Data collection
4.2.1.1 Selecting repositories
@@
We want to find repositories that are well covered by unit tests, so that we can automatically evaluate the migration correctness using the tests. We start by downloading a list of 95,472 Python GitHub repository metadata using SEART seart service. We use the filters to select Python repositories that have at least one commit after January 1, 2024 to discard repositories that are not actively maintained. We also filter out repositories with less than 100 lines of code to ensure that the repositories have some code to migrate.
For each repository, SEART metadata includes many details, including the repository name, the latest commit, the lines of code in the repository and repository size in KB. Before cloning the repository, we use some heuristics to filter out the repositories that may not be suitable for evaluation.
@@
    No requirements files: We use requirements files to build the virtual environment. We filter out 39,916 repositories that do not have a requirements.txt file in the root directory, leaving us with 24,993 repositories. We check this by making an HTTP head request to the GitHub raw file URL. If the response code is not 200, we conclude that the file does not exist.
@@
    Filtering by test files: Python test file names usually contain the term test. We use GitHub code search API [^4] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contains the term test, and filter out a repository if it does not have any such files. We filter out 12,558 repositories that do not have any test files, leaving us with 12,435 repositories.
@@
    Could not run tests: We attempt to setup a virtual environment for the repository using the requirements file and run the tests following Section 3.3.2. Due to various repository specific requirements; e.g., python version, system requirements, incomplete requirements file; we are often unable to run the tests. We filter out 10,104 repositories that we could not run the tests for, leaving us with 2,314 repositories.
@@
    Tests require user interaction: Some repositories require user interaction for the test to continue. For example, one repository requires a user to select an application to run a file. We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation. We store the commit SHA for the version of the repository we use to ensure reproducibility.
Selecting library pairs We collect the dependencies of the 798 selected repositories by parsing their requirements files. These dependent libraries are the potential source libraries. Then for each source library, we find a list of candidate analogous libraries using GPT 4o mini, and then manually we manually verify them. We describe the process below.
Collecting source libraries For each repository, we parse the requirements file to find the dependencies. This gives us an initial list of 1,421 unique source libraries. We then look up the library metadata on libraries.io libraries.io . We exclude 12 libraries that are not found on libraries.io, leaving us with 1,409 libraries. To ensure that the libraries are commonly known, we filter out 602 libraries that have less than 100 dependents, and filter out 2 libraries that do not have a description, leaving us with 805 potential source libraries.
@@
1Identify alternative Python libraries that offer similar functionality as <lib_name>,
2allowing for potential replacement in applications.
@@
3The alternative libraries should be available on PyPI.
@@
4Produce the result as a JSON list, having only the library names.
5The produced library names should match the PyPI package names.
@@
Finding analogous library pairs For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3. This gives us a total of 4,856 candidate library pairs. All of these pairs may not actually be analogous, i.e., they may not be suitable for migration, so we need to verify them. This manual task can be very time consuming, therefore, we first automatically filter out some of the pairs. First, we filter out 4,215 pairs where the source library is used in less than 10 of the selected repositories. We do this to ensure that every included library pair has at least 10 repositories to evaluate on, so that we can get a good estimate of the performance. Then, we filter out 192 pairs with unmaintained potential target library. Specifically, we exclude the target libraries that have not been updated since January 1, 2024. This leaves us with 449 pairs to manually verify.
During manual evaluation, we look into the library documents and confirm 60 library pairs that are indeed analogous and can be potentially migrated. We exclude the remaining because they fall into one of the below categories:
  1. i.
    Not analogous: Some libraries are not analogous, e.g., they provide different functionality or have different APIs.
  2. ii.
    Test related libraries: For example, pytest, mock, etc. We exclude these libraries because our evaluation process relies on the tests. Therefore, migrating the test libraries themselves would may break our automated evaluation.
  3. iii.
    Non-API tools: Some dependencies are command line interfaces. Since they are not used in the code, we ignore them. For example, black, flake8, etc.
  4. iv.
    Documentation tools: Some dependencies are used for documentation generation. Since they are not used in the code, we ignore them. For example, sphinx, mkdocs, etc.
4.2.1.2 Collecting migrations
From the above process, we have a list of 798 repositories and 60 analogous library pairs. For each of these repositories, we now find which of the source libraries from the analogous library pairs are used in the repository. To find actual usage, not just listing in the requirements file, we parse the source code of the repository and look for import statements that use the source libraries. We find that 175 repositories use at least one of the source libraries. Each source library can have one or more analogous target libraries. A repository, a commit, a source library and a target library thus form a unique migration instance. We get a total of 717 unique migration instances, which is the product of 175 repositories and 48 library pairs. These numbers are less than the 798 repositories and 60 analogous library pairs we started with, because we find many instances where a repository declares a dependency on a library, but does not actually use it in the code.
To summarize, our evaluation is based on 717 migrations between 48 library pairs in 175 repositories. The 48 library pairs include 24 unique source libraries and 37 unique target libraries, and 55 unique libraries in total.
4.2.2 Running migrations
@@
In our initial LLM experiment (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run the migrations in this evaluation.
4.2.3 RQ-B1: How many migrations can PyMigTool perform correctly?
4.2.3.1 RQ-B1: Approach
@@
We describe in Section 4.1 that PyMigTool generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration round, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation. For a given post migration test report, we define correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the test report before the migration.
Table 6: Example of migration correctness calculation.
Test case TestReportpremig TestReportllmmig TestReportmerge TestReportasync
Test 1 Pass Pass Pass Pass
Test 2 Pass Fail Fail Fail
Test 3 Pass Fail Fail Pass
Test 4 Pass Fail Pass Pass
Test 5 Fail - - -
# passing 4 1 2 3
correctness 25% 50% 75%
We use Table 6 to explain our correctness calculation. Assume that there are fives tests in the project. TestReportpremig has 4 passing tests and 1 failing test. We only take into account the tests that passed in TestReportpremig, because the failing tests cannot cannot indicate whether the migration is correct or not. In TestReportllmmig, only 1 test passes out of 4, therefore, the correctness is 25%. Similarly, in TestReportmerge and TestReportasync, 2 and 3 tests pass respectively, therefore, the correctness is 50% and 75% respectively.
4.2.3.2 RQ-B1: Results
@@
The solid line indicate the median correctness, and the dashed line indicates the mean correctness.
@@
Table 7 summarizes our results, while comparing the different migration steps. The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations where all tests that passed before migration all pass after this migration step. Median correctness is the median correctness value across all evaluated migrations, where correctness is calculated according to Table 6. Number of migrations applied to is the number of migrations where this step was applicable to and got applied. For example, not all migrations require merging unchanged code or applying async transformations. Number of migrations improved shows the number of migrations where correctness improved in comparison to the previous step. Average improvement in correctness is calculated for those migrations where correctness improved. Figure 9 also shows the distribution of correctness of migrations after the each migration step.
All three violins have concentrated values at 0% and 100%, indicating that most of the migrations either pass all tests or fail all tests. From this figure and Table 7, we see that the median correctness of migration increased from 65% in the LLM migration round to 80% after the merge skipped code, and further increased to 86% after the infer async transformation. The distributions show that after post-processing, there are fewer migration with zero or low correctness, and more migrations with high correctness. This indicates that the post-processing steps are effective in improving the correctness of migrations.
@@
Note that the post processing steps are applied only when the migrations are not fully correct, and when that specific post processing step is applicable. For example, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration. On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
4.2.4 RQ-B2: How much manual effort is required to complete the incomplete migrations?
In RQ-B1, we see that PyMigTool is able to fully migrate 32% of the migrations; however, this means that 68% of the migrations would still require the developer to perform edits or fixes to complete the migrations. In this RQ, we look into how much additional manual effort is required to correct the incomplete migrations.
4.2.4.1 RQ-B2: Approach
@@
RQ-B1 shows that 233 migrations are completely correct after the automated migration process. The remaining 484 migrations will require fixes for the tests to pass. These 484 migrations are between 46 unique library pairs in 147 repositories. Attempting to fix all these migrations manually would be very time consuming and impractical. Therefore, we select one random migration from each library pair, while maximizing the number of unique repositories. This gives us a total of 43 migrations between 43 library pairs in 39 unique repositories.
For each of these migrations, we open the code using the PyCharm IDE at the version after PyMigTool migration. We refer to library documents and other online resources to understand each test error and how to fix it. After applying the edits, we runs the tests again and save the test report, which we refer to as TestReportmanual.
We estimate the development effort required for a migration by measuring the number of lines changed for the migration. For a migration, we compute a diff between the code before manual edit (i.e., the code after automating migration by PyMigTool) and the code after manual edits. From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines. We refer to this number as $MigLOC_{manual}$. Now, while this number gives us the absolute number of line changed, it does not give us the relative effort in comparison to the automated migration by PyMigTool. For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines. Therefore, we also measure the number of lines changed by PyMigTool during the automated migration from the diff between the code before automatic migration and the code after migration by PyMigTool. We refer to this number as $MigLOC_{auto}$. From these two numbers, we calculate the manual migration effort as the ratio of manual changes to the total changes.
$\mathit{Effort}_{\mathit{manual}}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (1)
Similarly, automatic migration effort is calculated as the ratio of automatic changes to the total changes.
$\mathit{Effort}_{\mathit{auto}}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}$ (2)
Note that the total migration effort is the sum of $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$, i.e., $\mathit{Effort}_{\mathit{manual}}+\mathit{Effort}_{\mathit{auto}}=1$.
4.2.4.2 RQ-B2: Results
@@
Figure 10: Distribution of effort savings across manually fixed migrations.
The yellow dotted line shows median effort saving across all 24 migrations.
We apply manual edits to the selected 43 migrations. The manual edit effort fixes at least one failing test for 24 (56%) migrations, including 21 (49%) migrations that were completely fixed.
@@
For the 21 migrations that we were able to fully correct through the manual edits, we measure $\mathit{Effort}_{\mathit{manual}}$and $\mathit{Effort}_{\mathit{auto}}$using Equation 1 and Equation 2, respectively. Figure 10 shows the proportion of automatic and manual changes made for each migration. The blue part of the stacked bar is the $\mathit{Effort}_{\mathit{auto}}$, while red part is $\mathit{Effort}_{\mathit{manual}}$. The yellow dotted line shows the median of $\mathit{Effort}_{\mathit{auto}}$  across all 21 migrations, which is 86%. This means that for a median migration, PyMigTool performed 86% of the changes, while the developers only needed to perform 14% of the changes.
4.3 Implications and discussion
We now discuss some of the issues we noticed with the LLM-based migration in PyMigTool and potential extensions of PyMigTool to resolve these issues in the section after this (Section 4.3.2).
4.3.1 Reoccurring issues in PyMigTool migrations
To understand current limitations of PyMigTool, we look at the manual fixes we had to perform. The examples we show here are from the original migrations from our experiments, but often adopted for clarity and conciseness. In the left-to-right or top-to-bottom comparisons, the changes are highlighted in red.
Original code
1class EncAlchemyField:
2 impl = BINARY
3
4 def __init__(self, key):
5 self.key = key
6 self.crypto=Crypto(key)
PyMigTool migration
1class EncTortoiseField:
2 SQL_TYPE = "BLOB"
3
4 def __init__(self, key):
5 self.key=key
6 self.crypto=Crypto(key)
Manual fix
1class EncAlchemyField:
2 SQL_TYPE = "BLOB"
3
4 def __init__(self, key):
5 self.key=key
6 self.crypto=Crypto(key)
Figure 11: Example of PyMigTool renaming a public class, and we manually fixing it back to original name in a sqlalchemy→tortoise-orm migration.
  1. Issue-1
@@
    The LLM refactors public program elements: In the LLM migration round, PyMigTool migrates each file in isolation, which can lead to incorrect migrations. For example, in Figure 11, PyMigTool renames the public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name. This class is used in other parts of the code. Renaming the file prevents its callers from finding it, causing the errors in tests. Since PyMigTool does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class. Note that the prompt (2) does ask not to change any API names, however, the LLMs do not always strictly follow the instructions.
@@
    5def test_get_env_vars_optional(self):
    6 expected_result = EnvVars(
@@
    11 expected = str(expected_result)
    12 self.assertEqual(actual, expected)

### v2

@@ 4.4 Post-processing Step 2: Infer async transformation
3def fetch_data(year):
@@
3def test_fetch_data():
@@
9
10.
After Infer async transformation
@@
3async def fetch_data(year):
@@
Identify asynced functions
We call a function asynced if it was synchronous before migration, but is asynchronous after the LLM migration. For a given migrated file, MigrateLib creates an abstract syntax tree (AST) of the code before migration and another AST after migration. It then finds the function definitions that are not asynchronous in the AST before migration, but are asynchronous in the AST after migration. Since the location (line number) of the function definition may change due to the migration, MigrateLib matches the functions by their qualified name. In Figure 7, the function fetch_data is asynced.
Identify to-async functions
@@
If a function foo calls a function bar, and bar calls a function fizz, then foo is said to transitively call bar and fizz. For a given asynced function, MigrateLib finds all its transitive callers using the call graph, and marks them as to-async, meaning that the async keyword must be added to their function definition. For example, in Figure 7, the function fetch_data in service.py is asynced after the LLM migration. This function is called by the function test_fetch_data in service_test.py, therefore, MigrateLib marks test_fetch_data as to-async.
Identify to-await calls
MigrateLib again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await, meaning that an await keyword must be added to these calls. For example, in the call to fetch_data in the function test_fetch_data (line 5, after LLM migration) is marked as to-await, because it is a call to an asynced function.
Upon finding all asynced functions, to-async functions, and to-await calls, MigrateLib applies the necessary code transformations using the LibCST library contributors (2025). Additionally, MigrateLib decorates any asynced and to-async test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests. After finishing the async transformations, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportasync.
5 Phase 3: Evaluating MigrateLib
In this phase, we want to evaluate MigrateLib on a large set of real-world projects. To enable such large-scale evaluation, we rely on tests to determine if the migration is successful. Since the majority of repositories in PyMigBench did not contain tests, we resort to the idea of simulated migrations where we find source libraries used in a project and migrate the project to use an analogous target library.
Our evaluation aims to answer the following research questions:
  1. RQ-2.1
@@ 5 Phase 3: Evaluating MigrateLib
    How many migrations can MigrateLib perform correctly? We measure how well MigrateLib performs on the simulated migrations without any manual intervention.
  2. RQ-2.2
@@
    How much manual effort is required to complete the incomplete migrations? While the best case scenario is that MigrateLib can migrate the project without any issues, a more practical expectation is that MigrateLib would do most of the task correctly, but there are some parts that require manual intervention.
5.1 Data collection and selected model
@@
We want to find repositories that use some source library, and then migrate them to use an analogous target library. To make the test-based evaluation reliable, we want to select repositories whose tests have good code coverage. We start by downloading a list of 95,472 Python GitHub repository metadata using SEART Dabic et al. (2021) service. We use the filters to select Python repositories that have at least one commit after January 1, 2024 to discard repositories that are not actively maintained. We also filter out repositories with less than 100 lines of code to ensure that the repositories have some code to migrate. We then filter out repositories using the following criteria before cloning them for further analysis:
@@
    No requirements files: We use requirements files to build the virtual environment to be able to run the tests. We filter out 39,916 repositories that do not have a requirements.txt file in the root directory, leaving us with 24,993 repositories.
@@
    No test files: Python test file names usually contain the term test. We use GitHub code search API [^1] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contain the term test, and filter out a repository if it does not have any such files. We filter out 12,558 repositories that do not have any test files, leaving us with 12,435 repositories.
@@
    Could not run tests: We attempt to set up a virtual environment for the repository using the requirements file and run the tests following Section 3.1.3 (Evaluating LLM migrations using unit tests). Due to various repository-specific requirements, e.g., Python version, system requirements, incomplete requirements file, we are often unable to run the tests. We filter out 10,104 repositories that we could not run the tests for, leaving us with 2,314 repositories.
@@
    Tests require user interaction: We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation. We store the commit SHA for the version of the repository we use to ensure reproducibility.
We then collect the dependencies of the 798 selected repositories by parsing their requirements files. This gives us an initial list of 1,421 unique source libraries. We exclude 12 libraries that are not found on libraries.io Katz (2020), leaving us with 1,409 libraries. To ensure that the libraries are commonly known, we filter out 602 libraries that have less than 100 dependents, and filter out 2 libraries that do not have a description, leaving us with 805 potential source libraries.
@@
Identify alternative Python libraries that offer similar functionality as
@@
<lib_name>, allowing for potential replacement in applications. The alternative libraries should be available on PyPI.
@@
Produce the result as a JSON list, having only the library names. The produced library names should match the PyPI package names.
@@
For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3. This gives us a total of 4,856 candidate library pairs. All of these pairs may not actually be analogous, i.e., they may not be suitable for migration, so we need to verify them. This manual task can be very time-consuming; therefore, we first automatically filter out some of the pairs. First, we filter out 4,215 pairs where the source library is used in less than 10 of the selected repositories. This help us to keep the manual effort of evaluating analogous pairs manageable by only selecting pairs that appear in multiple projects in our data set. Then, we filter out 192 pairs with an unmaintained potential target library. Specifically, we exclude the target libraries that have not been updated since January 1, 2024. This leaves us with 449 pairs to manually verify.
During manual validation, we look into the library documentation and confirm 60 library pairs that are indeed analogous and can be potentially migrated. We exclude the remaining because they fall into one of the following categories (i) Not analogous:Some libraries are not analogous, e.g., they provide different functionality or have different APIs. (ii) Test related libraries:For example, pytest, mock, etc. We exclude these libraries because our evaluation process relies on the tests. Therefore, migrating the test libraries themselves may break our automated evaluation. (iii) Non-API tools:Some dependencies are command line interfaces. Since they are not used in the code, we ignore them. For example, black, flake8, etc. (iv) Documentation tools:Some dependencies are used for documentation generation. Since they are not used in the code, we ignore them. For example, sphinx, mkdocs, etc.
Now that we have a list of 798 repositories and 60 analogous library pairs, we want to find which repositories actually use the source libraries in our list. It could be the case that some repositories list a requirement that is not actually used in the code. Accordingly, we parse the source code of each repository and look for import statements that use the source libraries. We find that 175 repositories use at least one of the source libraries. Note that each source library can have one or more analogous target libraries (e.g., requests→aiohttp and requests→httpx). We use the term unique migration instance to refer to the combination of a repository, a commit, a source library, and a target library. We use the latest commit of each repository at the time of our experiment (January 2025) for this purpose. We find a total of 717 unique migration instances, spanning 175 repositories and 48 library pairs that we use for our experiments. The 48 library pairs include 24 unique source libraries and 37 unique target libraries, and 55 unique libraries in total.
@@
In our initial LLM empirical study (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run MigrateLib in this evaluation.
5.2 Evaluation metrics
@@
Recall from Section 4 that MigrateLib generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration step, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation. For a given post-migration test report, we define the correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the TestReportpremig. For example, if 80 tests passed in TestReportpremig, and 60 of those tests passed in TestReportllmmig, then the correctness after the LLM migration step is 75% (60/80).
It is possible that some migrations may not be fully correct even after all the MigrateLib steps. We want to see how much migration work is still needed after MigrateLib migration to get a fully correct migration. Therefore, we randomly select one migration instance from each of the library pair, and attempt to manually fix them.
After fixing a migration as much as possible, we run the tests again and generate a new test report, which we refer to as TestReportmanual. We then compute the correctness of the manually fixed version. For the migrations which become fully correct, i.e., 100% correctness, we estimate the amount of manual work needed to fix them to answer RQ-2.2. We use line number of modified lines to estimate the amount of manual change required to complete a migration. For a migration, we compute a diff between the code before manual migration and the code after manual migration. From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines. We refer to this number as $\mathit{MigLOC}_{\mathit{manual}}$. Now, this number gives us the absolute number of lines changed, therefore we cannot directly compare this number across different migrations. For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines. Therefore, we also measure the number of lines changed by MigrateLib during the automated migration from the diff between the code before automatic migration and the code after migration by MigrateLib. We refer to this number as $\mathit{MigLOC}_{\mathit{auto}}$. Accordingly, the total number of lines that needed to be changed for the migration is the sum of these two numbers. We use these to normalize the $\mathit{MigLOC}_{\mathit{manual}}$ as follows:
$\mathit{MigLOC}_{\mathit{manual}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (1)
Similarly, normalized automatic migration changes is calculated as the ratio of automatic changes to the total changes.
$\mathit{MigLOC}_{\mathit{auto}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (2)
5.3 RQ-2.1: How many migrations can MigrateLib perform correctly?
@@
The dashed lines indicate the mean, the solid lines indicate the mean.
@@
Table 5 summarizes the results of RQ-2.1, while comparing the different migration steps. The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations with 100% correctness. Median correctness is the median of correctness value across all evaluated migrations. Number of migrations applied to is the number of migrations where this step was applicable and thus got applied. Number of migrations improved shows the number of migrations where correctness improved in comparison to the previous step. Average improvement in correctness is the mean of improvement for the migrations where correctness improved after a post-processing step. Note that the post-processing steps are applied only when the migrations are not fully correct and when that specific post-processing step is applicable. For example, as shown in Table 5, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration. On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
The frequency curve chart in Figure 8 shows the distribution of correctness of migrations after each migration step. All three distributions have concentrated values at 0% and 100%, indicating that most of the migrations either fail all tests or pass all tests. We see that the median correctness of migration increased from 65% in the LLM migration step to 80% after the merge skipped code, and further increased to 86% after the infer async transformation. Also note that the LLM-only has the highest value at 0% correctness, which decreases after each post-processing step. Similarly, the value at 100% correctness increases after each post-processing step. The distributions show that after post-processing, there are fewer migrations with zero or low correctness, and more migrations with high correctness. This indicates that the post-processing steps are effective in improving the correctness of migrations.
RQ-2.1 MigrateLib is able to fully migrate 32% of the migrations without any manual intervention. The median correctness of migrations improves from 65% with only the LLM migration step to 86% after the post-processing steps.
5.4 RQ-2.2: How much manual effort is required to complete the incomplete migrations?
@@
RQ-2.1 shows that 233 migrations are completely correct after the automated migration process. The remaining 484 migrations will therefore require the developers to fix/complete them. To mimic what developers would do, we attempt to fix the migrations ourselves. However, the 484 migrations are between 43 unique library pairs in 147 repositories. To reduce the amount of manual work needed, while keeping the dataset diverse, we attempt to manually fix one incomplete migration from each library pair. Our selection results in a total of 43 migrations between 43 library pairs in 39 unique repositories, which we proceed to fix.
The yellow dotted line shows median automatic change across all 24 migrations. The numbers on top of each bar show the total changed lines of code (automatic + manual) for that migration.
@@
Figure 9: Distribution of manual and automatic changes across manually fixed migrations.
We are able to fix 24 out of 43 migrations. We could not fix the remaining 19 due to one or more of the following reasons: (1) As outsiders to the project, some of the application logic was too complex for us to understand and fix the migration. (2) In one case, the LLM migrated to dataclasses instead of cattrs, which made the migration incorrect. Fixing this migration would require undoing the entire migration and re-migrating it to the correct library, which is beyond the scope of fixing an incomplete migration. (3) In some cases, the tests themselves need to be updated to account for the behavior of the new target library. In several cases, we could not figure the correct expectation in the test oracle (again due to the complexity of the application) or because the original tests had complex mocking of the source library APIs. (4) We did observe cases where additional dependent libraries require migration, beyond the source-target pair. We consider such migrations beyond scope of this work, thus we did not attempt to fix such migrations.
@@
For the 24 migrations we were able to completely fix, we measure $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$ and $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$ using Equation 1 and Equation 2, respectively. Figure 9 shows the proportion of automatic and manual changes made for each migration. The blue part of the stacked bar is the $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$, while the red part is $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$. The yellow dotted line shows the median of $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$  across all 24 migrations, which is 73%. This means that for a median migration, MigrateLib performed 73% of the changes, while the developers would only need to perform 27% of the changes. Note that 3 bars in the figure have no red part, this is because the manual fixes required to successfully complete the migration were not source code changes but some environment change, such as installing a missing dependency.
RQ-2.2 For the migrations that required manual fixes, MigrateLib performed a median of 73% of the migration changes, while a developers would only need to perform 27% of the changes to complete the migration.
To understand where MigrateLib fell short, we qualitatively analyze the manual fixes we had to perform. We categorize them into the following observations. We provide (adapted) examples from the actual migrations.
  • The LLM often incorrectly uses target APIs: Sometimes the target API call used by the LLM does not correctly replace the source API call. This includes incorrect import statements, incorrect arguments to API calls, missing API calls, incorrect exception types, incorrect use of async/await, missing non-dafault argument, and other API related mistakes. For example, in a click→typer migration, the LLM did not use a name for the Typer() call, which was required. The equivalent Click API does not require a name, so MigrateLib missed adding it.
  • The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, especially when the code pattern is repetitive or does not have source API usage. We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in MigrateLib to merge the skipped code with the migrated code (see Section 4.3). However, the heuristics used in Merge skipped code still fail to identify some skipped code so we had to manually add it back in the fix.
  • MigrateLib struggles with complex async migrations: Some sync to async migrations require complex changes in the code structure, such as changing class constructors to async methods, changing sync context managers to async context managers. The LLM often struggles to perform such complex changes correctly. Our infer async transformation in tool only performs simple transformations, such as adding async/await keywords and changing context managers. We manually fixed some of these complex async migrations.
@@
  • The LLM refactors public program elements: In the LLM migration step, MigrateLib migrates each file in isolation, which can lead to incorrect migrations. For example, In the sqlalchemy→tortoise-orm migration, MigrateLib renames a public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name. This class is used in other parts of the code. Renaming the file prevents its callers from finding it, causing the errors in tests. Since MigrateLib does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class. Note that the prompt (2) does ask not to change any API names, however, the LLMs do not always strictly follow the instructions.
@@
    5def test_get_env_vars(self):
    6 expected_var = EnvVars(
@@
    11 expected = str(expected_var)
    12 assert actual == expected